A basketball team sells tickets that cost $10, $20, or for VIP seats, $20. The team has sold 3320...

A basketball team sells tickets that cost $10, $20, or for VIP seats, $30. The team has sold 3320 tickets overall. It has 296 more $20 tickets than $10 tickets. The total sales are $65,050. How many tickets of each kind have been sold?
Answer #1

here let x,y and z are tickets sold for $10, $20 and for VIP seats category

hence x+y+z =3320 ...........(1)

10x+20y+30z =65050...(2)

y=296+x .............(3)

solving above three equations:

number of $10 tickets =1053

number of $20 tickets =1349

number of VIP tickets =918